hey,
please i have a question on one of my word problem,
what is the smallest possible total of a number and its square?
i will be waiting for the answer,
and thank you

Answer by stanbon(75887)

what is the smallest possible total of a number and its square?
i will be waiting for the answer,
------------------------
Total = x + x^2
You have a quadratic with a=1 and b = 1
The minimum occurs when x = -b/2a = -1/2 = -1/2
------------
The miniumum total is T(-1/2) = (-1/2) + (-1/2)^2 = (-1/2) + 1/4 = -1/4
==============================
Cheers,
Stan H.